Esther's age is one year more than twice Isabela's age. Eight years ago, the sum of their ages was equal to Esther's age 10 years hence. How old are they now?
Answer by ankor@dixie-net.com(22740)

e = Esthers age
i = Isabel's age
:
Write an equation for each statement.
:
Esther's age is one year more than twice Isabela's age.
e = 2i + 1
:
Eight years ago, the sum of their ages was equal to Esther's age 10 years hence.
(e-8) + (i-8) = e + 10
e + i - 16 = e + 10
e + i = e + 10 + 16
e + i = e + 26
subtract e from both sides; therefore
i = 26 yrs is Isabel's age
then
e = 2(26) + 1
e = 53 yrs is Esther's age
:
;
see if that checks out in the statement
"Eight years ago, the sum of their ages was equal to Esther's age 10 years hence."
45 + 18 = 53 + 10
63 = 63
:
How old are they now? 26 and 53